# Optimizing Instructions and Demonstrations for Multi-Stage Language Model Programs

### v1

@@ 6 Results & Discussion
Lesson 2: Optimizing both instructions and few-shot examples with MIPRO generally yields the best overall performance. We support this with a statistical test comparing MIPRO with the second highest averaging optimizer for each task. The two exceptions to this are HotPotQA and Heart Disease. We hypothesize that instructions are less valuable for tasks like HotPotQA, whose final module is a relatively straightforward Q&A task that is likely in-distribution for many models. For Heart Disease, we hypothesize that this is due to initializing our optimizers with a simple seed instruction that does not convey any classification criteria, which current instruction optimizers have a limited ability to infer.
Lesson 3: Instruction optimization is most important for tasks with conditional rules that are (i) not immediately obvious to the LM and (ii) not expressible via a limited number of few-shot examples. This hypothesis is supported by our Iris experiment—and primarily by our HotPotQA Conditional experiments, where we optimize over a seed instruction stating the rules of the task. For this task, we find that even 0-shot instruction optimization outperforms demonstration-only optimization. In these cases, especially if the task is complex, it’s important that we optimize over a seed prompt, as our optimizers are not yet able to infer all task rules. (This lesson is also reflected in the Heart Disease results, as discussed above.)
@@ 8 Conclusion
We formalize the problem of optimizing LM programs by searching over the possible prompts (§2). We identify two challenges of LM program optimization: (1) proposal generation and (2) credit assignment during optimization. We address these challenges using three proposal generation and three credit assignment strategies (§3) and explore a representative subset of the algorithms (§4) using new benchmark of diverse tasks (§5.1). Our findings show that optimizing few-shot demonstrations is very powerful, but instruction optimization can be essential for complex task specifications, and jointly optimizing both is highly effective.
@@
This work studies a set of optimizers under a fixed budget, but does not examine how optimization dynamics might differ across extremely low or high budget scenarios. As discussed in Section 6, doing so may reveal new insights about the trade-offs between different optimizers, such as those that learn to improve proposals overtime, versus those that optimize over existing proposals very efficiently. Additionally, a limitation of the optimizers proposed in this work is that they have limited ability to infer the rules governing complex tasks without a handwritten seed prompt. Some information may be inferred from grounding (for example, information about the dataset or examples of the task), but this may not be enough to extrapolate to learning a set of complex rules. We encourage future work to explore how optimizers might learn such task dynamics without any handwritten inputs.
@@ Appendix A Detailed Task Descriptions
HoVer (Jiang et al., 2020) contains claims that require many-hop search steps over Wikipedia to be fact-checked. We consider the sub-task of retrieving all required evidence. Our LM program performs a search on the claim, summarizes the results (LM call 1), adds that to context and proposes the next search query (LM call 2), performs a search, summarizes the results (LM call 3), proposes a search query based on both summaries (LM call 4), and finally searches once more. HoVer has gold labels for the documents to be retrieved. We measure Retrieval@21 with all top-10 retrieved documents across three hops. We restrict HoVer to 3-hop examples in order to study optimization within a uniform and difficult (> 2 hop) setting.
@@ B.4 Language Model Hyperparameters
. For ScoNe this is 200 tokens, for HoVer this is 600 tokens, and for all other tasks this is 150 tokens. Importantly we run Llama 3 without a chat template as we find this behaves better given DSPy’s autocomplete prompt style. For our proposer model, we use GPT-3.5 with temperature 0.7 and top_p=1.0 or GPT-4 with the same settings for ScoNe and HoVer.
@@ E.5 Iris
E.6 Heart Disease
@@ F.2 Single-module OPRO
The starting point for OPRO is a meta-prompt, which gives the state of the optimizer at each iteration. The meta-prompt consists of meta-instructions, task instructions with their scores from training data evaluations, and task exemplars. The model is asked to generate a new candidate task instruction that is different from the ones already included in the meta-prompt.111In the OPRO paper, a set of candidate instructions is generated and scored at each step. For simplicity, we consider only a single candidate per step. This is scored, and then the meta-prompt is updated with the (possibly) revised set of top-scoring task instructions.

### v2

@@ 6 Results & Discussion
| Optimizer | ScoNe | HotPotQA | HoVer | HotPotQA Cond. | Iris | Iris-Typo | Heart Disease | |||||||||||
| Train | Dev | Test | Train | Dev | Test | Train | Dev | Test | Train | Dev | Test | Train | Test | Train | Test | Train | Test | |
| Instructions only (0-shot) | ||||||||||||||||||
| N/A | 57.0 | 56.2 | 69.1 | 35.4 | 31.8 | 36.1 | 30.2 | 30.8 | 25.3 | 13.8 | 10.5 | 6 | 46.4 | 40.9 | 34.7 | 32 | 23.3 | 26.8 |
| Module-Level OPRO G | 70.0 | 67.4 | 76.1 | 36.0 | 31.7 | 36.0 | 30.0 | 30.0 | 25.7 | – | – | – | – | – | – | – | – | – |
| Module-Level OPRO | 69.1 | 67.6 | 73.5 | 41.9 | 36.2 | 39.0 | 37.1 | 38.6 | 32.5 | – | – | – | – | – | – | – | – | – |
| 0-Shot MIPRO | 66.3 | 65.2 | 71.5 | 40.2 | 34.2 | 36.8 | 37.7 | 38.4 | 33.1 | 22.6 | 20.3 | 14.6 | 40.8 | 36.4 | 56.8 | 56.7 | 26.8 | 25.8 |
| 0-Shot MIPRO++ | 69.0 | 66.9 | 75.7 | 41.5 | 36.2 | 39.3 | 37.1 | 37.3 | 32.6 | – | – | – | – | – | – | – | – | – |
| Demonstrations only (Few-shot) | ||||||||||||||||||
| Bootstrap RS | 74.9 | 69.6 | 75.4 | 48.6 | 44.0 | 45.8 | 42.0 | 42.0 | 37.2 | 16.4 | 15.0 | 10.4 | 95.2 | 94.1 | 58.9 | 58.7 | 78.4 | 79.2 |
| Bayesian Bootstrap | 75.4 | 67.4 | 77.4 | 49.2 | 44.8 | 46.2 | 44.6 | 44.7 | 37.6 | – | – | – | – | – | – | – | – | – |
| Both (Few-shot) | ||||||||||||||||||
| MIPRO | 74.6 | 69.8 | 79.4 | 49.0 | 43.9 | 46.4 | 44.7 | 46.7 | 39.0 | 28.4 | 28.1 | 23.3 | 98.4 | 88.6 | 69.1 | 68.7 | 75.2 | 74.2 |
@@
Lesson 2: Optimizing both instructions and few-shot examples with MIPRO generally yields the best overall performance. We support this with a statistical test comparing MIPRO with the second highest averaging optimizer for each task. The exceptions to this are HotPotQA, Heart Disease, and Iris without a typo. We hypothesize that instructions are less valuable for tasks like HotPotQA, whose final module is a relatively straightforward Q&A task that is likely in-distribution for many models. For Heart Disease, we hypothesize that this is due to initializing our optimizers with a simple seed instruction that does not convey any classification criteria, which current instruction optimizers have a limited ability to infer.
Lesson 3: Instruction optimization is most important for tasks with conditional rules that are (i) not immediately obvious to the LM and (ii) not expressible via a limited number of few-shot examples. This hypothesis is supported by our Iris-Typo experiment—and primarily by our HotPotQA Conditional experiments, where we optimize over a seed instruction stating the rules of the task. For this task, we find that even 0-shot instruction optimization outperforms demonstration-only optimization. In these cases, especially if the task is complex, it’s important that we optimize over a seed prompt, as our optimizers are not yet able to infer all task rules. (This lesson is also reflected in the Heart Disease results, as discussed above.) In the Iris-Typo setting, our instruction optimizer even helps correct mistakes in the seed prompt.
@@ 8 Conclusion
We formalize the problem of optimizing prompts in LM programs (§2). We identify two key challenges of LM program prompt optimization: (1) proposal of a small set of high-quality prompts and (2) credit assignment during optimization. We address these challenges using three proposal generation and three credit assignment strategies (§3) and explore a representative subset of optimizer algorithms (§4) using a new benchmark of diverse tasks (§5.1). Our findings show that optimizing few-shot demonstrations is very powerful, but instruction optimization can be essential for complex task specifications with multiple conditional rules. Finally, we find that jointly optimizing both instructions and demonstrations using the MIPRO optimizer is the most effective approach in five out of seven settings.
@@
This work studies a set of optimizers under a fixed budget, but does not examine how optimization dynamics might differ across extremely low or high budget scenarios. As discussed in Section 6, doing so may reveal new insights about the trade-offs between different optimizers, such as those that learn to improve proposals overtime, versus those that optimize over existing proposals very efficiently. In our experiments, we also use a fixed proposer LM and task LM. Future research should assess whether the proposed methods demonstrate consistent performance when employing different models. Furthermore, a limitation of the optimizers introduced in this work is their restricted ability to infer the rules governing complex tasks without a handwritten seed prompt. While some information can be gleaned from grounding—such as dataset details or examples of the task—this may be insufficient for extrapolating a comprehensive set of rules. We encourage subsequent studies to investigate how optimizers could learn such task dynamics without relying on handwritten inputs. Finally, while we have made efforts to establish a benchmark that covers a diverse range of tasks and programs, there remains more to learn regarding the performance of optimizer methods on increasingly complex tasks and programs. Improving this benchmark presents a promising avenue for future research.
Acknowledgements
Krista Opsahl-Ong is supported by the NSF CS-Grad4US Graduate Fellowship. This work was partially supported by IBM as a founding member of the Stanford Institute for Human-Centered Artificial Intelligence (HAI), and by the HAI Hoffman–Yee Grant “Dendritic Computation for Knowledge Systems”. Additional support was provided by the Wallenberg AI, Autonomous Systems and Software Program (WASP), funded by the Knut and Alice Wallenberg Foundation, as well as by Digital Futures at KTH. Finally, this research was also supported in part by affiliate members and other supporters of the Stanford DAWN project – Facebook, Google, and VMware.
@@ Appendix A Detailed Task Descriptions
HoVer (Jiang et al., 2020) contains claims that require many-hop search steps over Wikipedia to be fact-checked. We consider the sub-task of retrieving all required evidence. Our LM program performs a search on the claim, summarizes the results (LM call 1), adds that to context and proposes the next search query (LM call 2), performs a search, summarizes the results (LM call 3), proposes a search query based on both summaries (LM call 4), and finally searches once more. HoVer has gold labels for the documents to be retrieved. We measure Retrieval@21 with all top-10 retrieved documents across three hops. We restrict HoVer to examples with 3 supporting facts in order to study optimization in a more challenging setting. We note that different supporting facts can originate from the same document, so finding all required documents can sometimes be done in <3 hops. Approximately 67% (83%) of the training (test) sets are 3-hop queries, and the rest are 2-hop queries. In follow-up studies, we recommend filtering by examples that require 3-hops to study retrieval in a more uniform setting.
@@ B.4 Language Model Hyperparameters
. For ScoNe this is 200 tokens, for HoVer this is 600 tokens, and for all other tasks this is 150 tokens. Importantly we run Llama 3 without a chat template as we find this behaves better given DSPy’s autocomplete prompt style. For our proposer model, we use GPT-3.5 with temperature 0.7 and top_p=1.0 or GPT-4 with the same settings for ScoNe, HoVer, and Iris.
@@ E.5 Iris
E.6 Iris-Typo
E.7 Heart Disease
@@ F.2 Single-module OPRO
The starting point for OPRO is a meta-prompt, which gives the state of the optimizer at each iteration. The meta-prompt consists of meta-instructions, task instructions with their scores from training data evaluations, and task exemplars. The model is asked to generate a new candidate task instruction that is different from the ones already included in the meta-prompt.222In the OPRO paper, a set of candidate instructions is generated and scored at each step. For simplicity, we consider only a single candidate per step. This is scored, and then the meta-prompt is updated with the (possibly) revised set of top-scoring task instructions.